OpenStudy (anonymous):

b over 6- b-2 over 8 and that equals 8 what is b

OpenStudy (anonymous):

can either of you help me

OpenStudy (saifoo.khan):

i can help, but please write the ques.

OpenStudy (anonymous):

b over 6 as a fraction b\6 minus the fraction b-2 over 8 and both fractions =4

OpenStudy (anonymous):

I thought it equaled 8

OpenStudy (saifoo.khan):

OK, wait,.

OpenStudy (anonymous):

b\6 - b-2/8 = 4

OpenStudy (anonymous):

sorry that is the correct question

OpenStudy (anonymous):

how did you get that

OpenStudy (anonymous):

wait

OpenStudy (saifoo.khan):

"Online Class"

OpenStudy (anonymous):

It's not -6

OpenStudy (anonymous):

what is it

OpenStudy (anonymous):

I'm workin on it

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

b=90

OpenStudy (anonymous):

can you explain how you did that

OpenStudy (anonymous):

If the questions is: b/6 - (b-2)/8 = 4, then b = 90. Did I get the question right?

OpenStudy (anonymous):

b over 6 minus b minus 2 over 8 = 4

OpenStudy (anonymous):

Yeah, LA-MrC got it

OpenStudy (anonymous):

Multiply both sides by 48 first

OpenStudy (anonymous):

Or you can multiply both sides by 24.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then what

OpenStudy (anonymous):

24[ b/6 - (b-2)/8] =[ 4]24 4b - 3(b-2) = 96 4b -3b + 6 = 96 b + 6 + 96 b = 90